OpenStudy (anonymous):

2h-14=-5h show work please

OpenStudy (anonymous):

If you will try then it will become simple.. Do you know how ??

OpenStudy (anonymous):

no i forgot how too

OpenStudy (anonymous):

Add 5h both the sides and can you tell me what did you get ?

OpenStudy (anonymous):

If you are not getting what I am saying then you can freely tell me... Tell me can you add 5h both the sides or not ??

OpenStudy (anonymous):

i believe it's 2h-14

OpenStudy (anonymous):

Yeah your belief is right.. You are to solve here for h.. You are to find h here and for finding h your first step should be to add 5h both the sides first.. Simply tell me can you do that or not ?

OpenStudy (anonymous):

no you cant

OpenStudy (anonymous):

I can't ??? Here you have to find h or not??

OpenStudy (anonymous):

Are you seeing "=" sign there in between 2h - 14 and -5h or not??

OpenStudy (anonymous):

Can you tell me what this equal implies ??

OpenStudy (anonymous):

yes there is an = between 2h-14 and -5h

OpenStudy (anonymous):

This means term on left hand side is equal to the term on right hand side.. Right ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

This is just same as : Suppose if you have 2 apples and I have also 2 two apples.. Now if someone will give 5 apples each to us then we have 7 apples and again we are having same number of apples..

OpenStudy (anonymous):

Now if I add 5h both the sides, then this will not affect the equal sign there "=".. So you can safely add 5h to both the sides, this I am doing because we have to find h here: So: 2h - 14 + 5h = -5h + 5h Or : 7h - 14 = 0 Now add 14 both the sides: 7h - 14 + 14 = 0 + 14 7h = 14 Now divide by 7 both the sides to get h.. h = 14/7 h = 2 Gotta go... Sorry I could not help you more.. @hartnn will help now..

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

2h-14=-5h 2h + 5h = 14 (transpose -5h to the other side as well as -14 & that turns out to be positive) 7h = 14 h = 2 So h is 2 :D

hartnn (hartnn):

@lyserdoll h=2 is correct, did u get how ?

OpenStudy (anonymous):

yeah i got it now thanks